OpenStudy (anonymous):

how that each of the following sequences {ai} is monotonic(or has a monotonic tail) and is bounded, and therefore converges by the Bounded Monotonic Sequence Theorem.

OpenStudy (anonymous):

\[\frac{ i }{ i^2 +1 }\]

OpenStudy (anonymous):

Would it be decreasing?

OpenStudy (sirm3d):

prove that either \(a_i\ge a_{i+1}\) (monotonically decreasing) or \(a_i\le a_{i+1}\) (monotonically decreasing) for all positive integers \(i\). my bet is on the first case (decreasing). assume the inequality and show that the solution to the inequality is all positive integers \(i\).

OpenStudy (anonymous):

wait so I did it out and found out it was decreasing but like what does that mean? Does that show that it is monotonic?

OpenStudy (sirm3d):

because it is decreasing, is is bounded above by the first term.

OpenStudy (anonymous):

Oh okay thanks

OpenStudy (sirm3d):

oh, don't forget \[\frac{1}{2}\ge a_i>a_{i+1}>0\] so the sequence converges to zero.